OpenStudy (anonymous):

integrate sqrt(1-x^2-y^2)dy

ganeshie8 (ganeshie8):

that looks like part of double integral ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}\int\limits_{-\sqrt{1-x^2}}^{\sqrt{1-x^2}}\int\limits_{0}^{\sqrt{1-x^2-y^2}}dzdydx\]

ganeshie8 (ganeshie8):

\(\large z = 1-y^2\) is not a cylinder

ganeshie8 (ganeshie8):

I don't see how that triple integral represents the region in first octant bounded by the given planes... could u check once ? :)

OpenStudy (anonymous):

http://www.math.csusb.edu/faculty/hasan/252_Test_1_Material.pdf question sixteen b

ganeshie8 (ganeshie8):

its a sphere !

ganeshie8 (ganeshie8):

you're working question sixteen b, or sixteen a ?

OpenStudy (anonymous):

oh i am sorry i am working for hours.its my mistake i posted wrong question

OpenStudy (anonymous):

b

ganeshie8 (ganeshie8):

equation of sphere of radius \(1\) and centered at origin : \(\large x^2 + y^2 + z^2 = 1\)

ganeshie8 (ganeshie8):

\(\large z^2 = 1-x^2 - y^2\) \(\large z = \pm \sqrt{1-x^2 - y^2}\)

ganeshie8 (ganeshie8):

Upper hemi-sphere : \(\large z = + \sqrt{1-x^2 - y^2}\) Lower hemi-sphere : \(\large z = - \sqrt{1-x^2 - y^2}\)

ganeshie8 (ganeshie8):

The question is asking u to just setup the triple integral for volume of upper sphere

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

can u sketch the hemisphere ^